OpenStudy (anonymous):

A class contributed $3.50 in nickels and dimes to the Red Cross. In all there were 45 coins. How many were there of each kind? please help? :(

OpenStudy (anonymous):

put \(x\) as the number of nickles, and \(y\) as the number of dimes and solve \[x+y=45\] \[5x+10y=350\]

OpenStudy (anonymous):

or if you like, say \(x\) is the number of nickles, \(45-x\) is the number of dimes and solve \(5x+10(45-x)=350\) may be quicker

OpenStudy (anonymous):

you good from there?

OpenStudy (anonymous):

yeah i put the equation as 10x+5(45-x)=350..problem is i don't know what the first step to solve it =/

OpenStudy (anonymous):

remove parentheses using distributive law and get \(10x+225-5x=350\)

OpenStudy (anonymous):

then combine like terms to get \(5x+225=350\) and you are done in two steps

OpenStudy (anonymous):

got it from here?

OpenStudy (anonymous):

yeah i got it thanks!

OpenStudy (anonymous):

yw